OpenStudy (anonymous):

What is the y-intercept of the line that has a slope of -1/2 and passes through the point (2,3)

OpenStudy (anonymous):

ok so 3=-1/2 (2) +x 3=-1+x the y intercept is 4

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no prob :)